### v1

# Personalize Before Retrieve: LLM-based Personalized Query Expansion for User-Centric Retrieval

Abstract
Retrieval-Augmented Generation (RAG) critically depends on effective query expansion to retrieve relevant information. However, existing expansion methods adopt uniform strategies that overlook user-specific semantics, ignoring individual expression styles, preferences, and historical context. In practice, identical queries in text can express vastly different intentions across users. This representational rigidity limits the ability of current RAG systems to generalize effectively in personalized settings. Specifically, we identify two core challenges for personalization: 1) user expression styles are inherently diverse, making it difficult for standard expansions to preserve personalized intent. 2) user corpora induce heterogeneous semantic structures—varying in topical focus and lexical organization—which hinders the effective anchoring of expanded queries within the user’s corpora space. To address these challenges, we propose Personalize Before Retrieve (PBR), a framework that incorporates user-specific signals into query expansion prior to retrieval. PBR consists of two components: P-PRF, which generates stylistically aligned pseudo feedback using user history for simulating user expression style, and P-Anchor, which performs graph-based structure alignment over user corpora to capture its structure. Together, they produce personalized query representations tailored for retrieval. Experiments on two personalized benchmarks show that PBR consistently outperforms strong baselines, with up to 10% gains on PersonaBench across retrievers. Our findings demonstrate the value of modeling personalization before retrieval to close the semantic gap in user-adaptive RAG systems. Our code is available at https://github.com/Zhang-Yingyi/PBR-code.
1 Introduction
Figure 1: An example comparing generic and personalized query expansion.
Retrieval-Augmented Generation (RAG) has emerged as a pivotal paradigm for enhancing the capabilities of large language models (LLMs) (Lewis et al. 2020; Zhao et al. 2024; Zhang et al. 2025), leveraging a two-stage process: first retrieving relevant external corpus, then generating responses conditioned on the retrieved information (Cheng et al. 2023). The effectiveness of this paradigm hinges on the quality of query expansion by using LLM world knowledge (Culpepper et al. 2021; Song and Zheng 2024), which directly impacts the accuracy and relevance of the retrieval stage. Most current strategies generate pseudo answers (Gao et al. 2023) or candidate completions (Jia et al. 2023) using LLMs in a zero-shot or few-shot manner. These generated semantic representations, often combined with the original query, are used to retrieve information from a global external corpus.
Recently, with LLMs widely adopted as personalized assistants, many user queries are required to retrieve from individual histories and contexts, demanding retrieval mechanisms that go beyond generic corpora to leverage user-specific data. Existing query expansion strategies remain limited in their ability to capture user-specific semantics. In practice, the same textual query may convey vastly different intentions depending on a user’s preferences (Zou et al. 2023), background knowledge (Westhofen, Jung, and Neider 2025), or expression style (Neelakanteswara, Chaudhari, and Zamani 2024). As users exhibit diverse personas and contextual needs, it becomes increasingly important for query expansion methods to be personalized—capable of adapting to user-specific semantics within private corpora—so as to ensure accurate retrieval and enhance downstream generation.
As illustrated in Figure 1 (top), when two users submit the same query—“How would you describe my dietary preferences?”—standard expansion methods yield generic outputs that fail to account for individual differences. This results in mismatched retrieval: for example, the health-conscious user (User 1) successfully retrieves content related to “organic ingredients,” which aligns with their preferences. In contrast, the variety-seeking user (User 2), whose personal content emphasizes “trying new side dishes,” fails to retrieve relevant information, highlighting the limitations of non-personalized expansion strategies. Such failures stem from neglecting rich signals in user—such as style, intent patterns, and thematic focus—resulting in semantically misaligned expansions and degraded retrieval performance.
Motivated by these limitations, we aim to develop a personalized query expansion framework that adapts to individual user intent and expression as shown in Figure 1. However, introducing personalization into expansion brings two fundamental challenges: (1) User expression styles are inherently diverse. Users articulate intent using varied linguistic patterns—ranging from minimal prompts to elaborative reasoning—driven by personal habits, domain familiarity, or rhetorical preference (Neelakanteswara, Chaudhari, and Zamani 2024). These styles are often implicit and non-transferable, making it difficult to construct expansions that faithfully preserve user-specific semantics within a generalizable framework. (2) User corpora induce heterogeneous semantic structures. Users’ personal corpora often differ significantly in topical coverage, content organization, and linguistic granularity. This high degree of heterogeneity makes it difficult to locate reliable semantic regions related to the user query, and in turn, complicates the alignment of user-specific preferences. Without user-specific semantic anchors to guide expansion, queries are prone to drifting toward irrelevant regions, resulting in mismatched retrieval despite plausible surface semantics. These challenges raise a central question: How can we personalize query expansion—both in style and structure—before retrieval to align with individual user intent and corpus characteristics?
To address these challenges, we propose a Personalize Before Retrieve (PBR) framework that integrates user-specific signals into query understanding prior to retrieval. PBR adopts a two-stage design that combines expression-level expansion with structure-level alignment, enabling queries to reflect both personal style and corpus semantics. ❶ P-PRF: Personal Style-Aligned Pseudo Relevance Feedback. This module extracts linguistic patterns from user history to guide LLMs in generating pseudo-utterances and reasoning paths. These signals capture personalized expression style and hidden intent, often missed by generic expansion methods. ❷ P-Anchor: Personal Structure-Aligned Semantic Anchoring. We encode user history into a semantic graph that captures localized relevance patterns. Personalized PageRank is applied to identify representative anchor points within this structure to encode the whole user corpora or to express user general preference. The query is then guided toward these anchors to reflect the user’s semantic landscape. By jointly modeling stylistic variation and structural relevance, P-PRF and P-Anchor produce personalized query representations that enhance alignment between user intent and retrieval semantics in RAG systems. Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to propose a personalized query expansion framework for RAG systems that adapts to the user for improving retrieval.
  • We propose P-PRF, a style-aware expansion module that generates pseudo queries and reasoning paths conditioned on user history.
  • We introduce P-Anchor, a graph-based alignment module that grounds queries within personalized semantic spaces via local corpus structure.
  • Experiments on two real-world dialogue datasets show that PBR significantly improves personalized retrieval, validating the benefits of pre-retrieval personalization.
2 Problem Definition
We study personalized query expansion, which expands underspecified queries to align with both user expression style and corpus heterogeneity informed by past interactions.
Formally, let $q$ represent the user query, and let $H=\{h_{1},h_{2},\dots,h_{n}\}$ denote the user’s historical utterances, referred to as the user corpus, where each $h_{i}$ is a textual segment from prior conversations or interactions. The query is encoded as $\mathbf{q}=\phi(q)\in\mathbb{R}^{d}$, while the historical utterances are encoded into a set of corpus vectors $\mathcal{C}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{n}\}$ using a fixed encoder $\phi(\cdot)$:
$\mathbf{c}_{i}=\phi(h_{i})\in\mathbb{R}^{d},\quad\forall i=1,\dots,n.$ (1)
The goal is to construct a personalized query representation $\mathbf{q}^{*}\in\mathbb{R}^{d}$ that retrieves content from the user corpus $\mathcal{C}$ not only based on lexical-semantic similarity, but also aligned with user-specific latent intent. Formally, the retrieved set is defined as:
$\mathcal{R}(\mathbf{q}^{*})=\text{Top}_{k}\left(\left\{\text{sim}(\mathbf{q}^{*},\mathbf{c}_{k})\mid\mathbf{c}_{k}\in\mathcal{C}\right\}\right),$ (2)
where $\text{sim}(\cdot)$ denotes a similarity function (e.g., cosine similarity), and $\text{Top}_{k}(\cdot)$ returns the top-$k$ most relevant items.
The core challenge lies in constructing $\mathbf{q}^{*}$ from the observable query $\mathbf{q}$ and the latent user-specific context $\mathcal{C}$. We formulate a transformation function $f_{\Theta}(\mathbf{q},\mathcal{C})$ such that:
$\mathbf{q}^{*}=f_{\Theta}(\mathbf{q},\mathcal{C})=\mathbf{q}+\Delta_{\text{user}}(\mathbf{q},\mathcal{C}),$ (3)
where $\Delta_{\text{user}}(\mathbf{q},\mathcal{C})$ encodes personalized adjustments reflecting individual expression style, intent formulation, and structural semantics in the user corpus.
Figure 2: Overview of PBR, which consists of three components: P-PRF for stylistic expansion, P-Anchor for structural grounding, and PBR Fusion module for generating the final query representation.
3 PBR Framework
3.1 Framework Overview
As illustrated in Figure 2, the PBR framework enhances personalized retrieval by explicitly incorporating user-specific signals into query representation before retrieval. It consists of three core modules: P-PRF simulates two complementary forms of feedback to generate pseudo utterances and pseudo reasoning that reflect the user’s stylistic tendencies and underlying reasoning logic. P-Anchor constructs a semantic graph over the user’s corpus and applies PageRank to identify representative anchor points that capture the corpus-level user anchor. These two sources of personalized signals are integrated via PBR Fusion, which computes a personalized and query-specific dynamic weight to balance the contributions of pseudo utterances and pseudo reasoning. The resulting fused representation $q^{*}$ reflects both user hidden style and contextual grounding, enabling accurate retrieval from the semantic index constructed over the user corpus $C$.
3.2 P-PRF
Generic query expansion methods often overlook the nuanced expression styles (Neelakanteswara, Chaudhari, and Zamani 2024) and implicit reasoning patterns behind user expression (Li et al. 2025), leading to mismatches in personalized retrieval. This challenge is exacerbated when queries are ambiguous or under-specified, requiring systems to infer intent beyond surface forms. Existing solutions typically apply uniform expansion strategies, failing to adapt to user-specific language signals.
To address these limitations, P-PRF simulates user-specific signals before retrieval by simulating personalized feedback in two complementary forms: roughly via stylistic utterances, and logically via intent reasoning. Rather than using the full user history $H$, we retrieve a task-relevant subset $H_{q}$ based on semantic similarity:
$H_{q}=\text{TopK}\big(\{\text{sim}(\mathbf{q},\mathbf{c}_{t})\}_{t=1}^{n}\big),\quad|H_{q}|=k_{1},$ (4)
where $\mathbf{q}$ and $\mathbf{c}_{t}$ are the embeddings of the query and past utterances. This subset captures salient contextual-aware user traits for downstream simulation.
Pseudo Utterance Generation (Roughly)
To simulate user-specific expression patterns, we employ a large language model $\mathcal{G}^{\text{utt}}_{\text{LLM}}$ to generate $m$ pseudo utterances conditioned on the original query $q$ and user dialogue history $H_{q}$. Each utterance $f_{i}$ is designed to reflect how the user might naturally articulate the query, capturing stylistic nuances such as tone, verbosity, and phrasing:
$F=\{f_{1},\dots,f_{m}\}=\mathcal{G}^{\text{utt}}_{\text{LLM}}(q,H_{q}),\quad\mathbf{f}_{\text{avg}}=\frac{1}{m}\sum_{i=1}^{m}\phi(f_{i}),$ (5)
where $\phi(f_{i})$ denotes the embedding of the $i$-th utterance and $\mathbf{f}_{\text{avg}}$ is their mean representation.
Pseudo Reasoning Generation (Logically)
Beyond surface expression, effective personalization also requires modeling the user’s implicit reasoning process. We define $\mathcal{G}^{\text{rea}}_{\text{LLM}}$ as a parallel simulating pipeline to elicit a step-by-step rationale $r$ of user expression:
$r=\mathcal{G}^{\text{rea}}_{\text{LLM}}(q,H_{q}),\quad\mathbf{r}=\phi(r).$ (6)
This rationale introduces logical cues that are often absent from surface queries, thereby providing a complementary semantic signal to guide expansion alignment.
Method Base HyDE Query2Term MILL CoT ThinkQE PBR (Ours)
Metrics R@5 N@5 R@5 N@5 R@5 N@5 R@5 N@5 R@5 N@5 R@5 N@5 R@5 N@5
[…]
Overall 0.4484 0.3669 0.3464 0.2945 0.3584 0.3060 0.3200 0.3254 0.3627 0.3000 0.4791 0.3902 0.5035 0.4201
Basic information 0.4515 0.3088 0.3106 0.2395 0.3424 0.2614 0.2879 0.2337 0.3424 0.2454 0.4606 0.3265 0.5091 0.3647
Preference (hard) 0.3659 0.3759 0.3122 0.3079 0.3659 0.3491 0.2927 0.4250 0.3171 0.3335 0.4195 0.4310 0.4049 0.4175
Social 0.4852 0.4356 0.3909 0.3259 0.3494 0.3144 0.3808 0.3923 0.4009 0.3320 0.5503 0.4554 0.5541 0.4914
Preference (easy) 0.4904 0.4582 0.4615 0.4422 0.4327 0.4095 0.3750 0.4195 0.4423 0.4133 0.5064 0.4626 0.5321 0.5129
[…]
Overall 0.3783 0.3074 0.3747 0.3144 0.3908 0.3256 0.3030 0.3059 0.3721 0.3098 0.3861 0.3393 0.4516 0.3855
Basic information 0.3515 0.2644 0.4015 0.2975 0.3894 0.2914 0.3030 0.2547 0.3455 0.2696 0.3455 0.2677 0.4515 0.3485
Preference (hard) 0.4000 0.3547 0.3512 0.3561 0.3805 0.3798 0.3220 0.4171 0.3463 0.3490 0.3366 0.3486 0.4341 0.4352
Social 0.3921 0.3048 0.2805 0.2435 0.3984 0.3163 0.2638 0.2829 0.4160 0.3046 0.4780 0.4334 0.4494 0.3777
Preference (easy) 0.4295 0.4199 0.4904 0.4646 0.3974 0.4038 0.3526 0.3944 0.4359 0.4285 0.4487 0.4357 0.4840 0.4800
[…]
Overall 0.3738 0.3015 0.3108 0.2597 0.3007 0.2497 0.2791 0.2869 0.3199 0.2642 0.3643 0.3156 0.4029 0.3402
Basic information 0.3970 0.2748 0.2955 0.2051 0.3000 0.1955 0.2879 0.2319 0.3152 0.2052 0.3152 0.2430 0.4121 0.3057
Preference (hard) 0.3268 0.3343 0.3073 0.3148 0.2976 0.3186 0.2244 0.3566 0.3073 0.3281 0.3463 0.3635 0.3707 0.3889
Social 0.3204 0.2799 0.2714 0.2443 0.2852 0.2503 0.2752 0.3004 0.2965 0.2647 0.4261 0.3735 0.3657 0.3089
Preference (easy) 0.4583 0.4065 0.4615 0.4356 0.3397 0.3693 0.3365 0.3819 0.4071 0.4117 0.4744 0.4289 0.4904 0.4734
Overall Average 0.4002 0.3253 0.3440 0.2895 0.3500 0.2938 0.3007 0.3061 0.3295 0.3121 0.4098 0.3484 $\textbf{0.4527}^{*}$ $\textbf{0.3819}^{*}$
Table 1: Retrieval performance in PersonaBench. The best results are in bold, and the second-best results are underlined. “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
3.3 P-Anchor
Existing methods are often designed to retrieve from a unified RAG corpus, lacking explicit grounding in the structural organization of individual user corpora (Tan et al. 2025; Wu et al. 2025). To address this, we propose P-Anchor, a structure-aware module that captures corpus-level preferences by identifying semantically central regions within a graph-structured user space.
Graph Construction from User Corpus.
We represent the user’s prior corpus—such as interaction records with AI assistants—as a semantic graph $G=(\mathcal{C},\mathcal{E})$, following previous study (Tang et al. 2025), where $\mathcal{C}$ and $\mathcal{E}$ denote the sets of nodes and edges, respectively. The set of nodes $\mathcal{C}$ consists of history corpus vectors, where each node $\mathbf{c}_{i}$ as in Eq. (1) represents an encoded segment. The edges $\mathcal{E}$ are established based on pairwise similarity between the nodes. For each pair, we compute cosine similarity:
$S_{ij}=\frac{\langle\mathbf{c}_{i},\mathbf{c}_{j}\rangle}{\|\mathbf{c}_{i}\|\cdot\|\mathbf{c}_{j}\|}.$ (7)
We construct a sparse adjacency matrix $A$ to represent tht structural links within the user corpus, which retains only the top-$k_{2}$ neighbors exceeding a similarity threshold $\theta$:
$A_{ij}=\begin{cases}S_{ij},&\text{if }S_{ij}\geq\theta\text{ and }j\in\text{top-}k_{2}(i),\\ 0,&\text{otherwise}.\end{cases}$ (8)
Graph-Based Anchor Representation.
We following PageRank (Page et al. 1999; Haveliwala 2002) over $G$ to estimate a stationary distribution $\boldsymbol{\pi}$ reflecting node centrality. Let $\mathbf{P}=norm(\mathbf{A})$ be the row-normalized transition matrix, then:
$\boldsymbol{\pi}^{(t+1)}=\alpha\cdot\mathbf{P}^{\top}\boldsymbol{\pi}^{(t)}+(1-\alpha)\cdot\frac{1}{n}\mathbf{1}.$ (9)
The user anchor is computed as a weighted aggregation:
$\mathbf{c}_{\text{Anchor}}=\sum_{i=1}^{n}\boldsymbol{\pi}_{i}\cdot\mathbf{c}_{i}.$ (10)
This representation encodes structurally central semantics from the user’s corpus, anchoring the query in a context-aware and personalized semantic space.
3.4 PBR Fusion
To synthesize the signals captured in P-PRF and P-Anchor, we design a fusion mechanism that integrates stylistic pseudo-feedback with the query and user anchor. Recognizing that the utility of pseudo-feedback varies across users and queries—depending on individual expression patterns and the semantic openness of the query—we introduce a personalized and query-specific dynamic weight that balances the contribution of P-PRF according to its semantic proximity to both the original query $\mathbf{q}$ and the corpus anchor $\mathbf{c}_{\text{graph}}$. Specifically, we compute two weights $w_{1}$ and $w_{2}$ to quantify the semantic alignment of the pseudo-utterance and reasoning feedback, respectively:
$w_{1}=1+\text{sim}\left(\frac{\mathbf{q}+\mathbf{c}_{\text{Anchor}}}{2},\mathbf{f}_{\text{avg}}\right),\quad$ (11)
$w_{2}=1+\text{sim}\left(\frac{\mathbf{q}+\mathbf{c}_{\text{Anchor}}}{2},\mathbf{r}\right),$ (12)
where $\text{sim}(\cdot)$ denotes cosine similarity. The additive constant ensures positivity and smooth interpolation.
Then, we define a user-specific personalization shift as:
$\Delta_{\text{user}}(\mathbf{q},\mathcal{C})=\mathbf{c}_{\text{Anchor}}+w_{1}\cdot\mathbf{f}_{\text{avg}}+w_{2}\cdot\mathbf{r},$ (13)
and construct the final personalized query embedding as:
$\mathbf{q}^{*}=\mathbf{q}+\Delta_{\text{user}}(\mathbf{q},\mathcal{C}).$ (14)
The resulting query $\mathbf{q}^{*}$ incorporates user-specific semantics from both style and structure. It encodes not only the immediate surface form of the input query, but also latent stylistic preferences, goal-driven reasoning, and historically salient corpus signals.
Final Retrieval.
We retrieve from the user corpus $\mathcal{C}$ using faiss-based (Douze et al. 2024) nearest neighbor search with the fused query vector $\mathbf{q}^{*}$. This process leverages embedding-level alignment to recover user-relevant content.
Dataset LongMenEval-s LongMenEval-m
Metrics R@1 N@1 R@3 N@3 R@5 N@5 R@1 N@1 R@3 N@3 R@5 N@5
[…]
Base 0.1885 0.7924 0.6730 0.8020 0.8067 0.8301 0.1098 0.5251 0.4057 0.5559 0.5394 0.6038
HyDE 0.1957 0.7852 0.6945 0.8173 0.8115 0.8436 0.1146 0.4964 0.4129 0.5411 0.5179 0.5828
Query2Term 0.1814 0.7446 0.6396 0.7684 0.7470 0.7959 0.1146 0.4749 0.3747 0.5091 0.4821 0.5531
MILL 0.1838 0.7709 0.6635 0.7845 0.6635 0.7712 0.1122 0.5155 0.4129 0.5552 0.4129 0.5468
CoT 0.1862 0.7470 0.6516 0.7707 0.7542 0.8006 0.1098 0.4654 0.3866 0.5110 0.4726 0.5475
ThinkQE 0.1933 0.8019 0.7064 0.8214 0.8162 0.8424 0.1169 0.5227 0.4168 0.5647 0.5517 0.6156
PBR $\mathbf{0.2100^{*}}$ $\mathbf{0.8210^{*}}$ $\mathbf{0.7279^{*}}$ $\mathbf{0.8444^{*}}$ $\mathbf{0.8282^{*}}$ $\mathbf{0.8598^{*}}$ $\mathbf{0.1217^{*}}$ $\mathbf{0.5537^{*}}$ $\mathbf{0.4320^{*}}$ $\mathbf{0.5776^{*}}$ $\mathbf{0.5537^{*}}$ $\mathbf{0.6177^{*}}$
[…]
Base 0.2196 0.8234 0.7399 0.8477 0.8568 0.8731 0.1265 0.5322 0.4415 0.5857 0.5346 0.6191
HyDE 0.2172 0.8138 0.7208 0.8372 0.8377 0.8599 0.1074 0.4988 0.4081 0.5483 0.5131 0.5884
Query2Term 0.2196 0.8019 0.6778 0.8000 0.8138 0.8351 0.1193 0.4940 0.3747 0.5209 0.4964 0.5689
MILL 0.2100 0.8162 0.7208 0.8382 0.7208 0.8257 0.1217 0.5179 0.3962 0.5530 0.3962 0.5450
CoT 0.2124 0.7709 0.6730 0.7932 0.7780 0.8187 0.1193 0.4487 0.3389 0.4920 0.4344 0.5273
ThinkQE 0.2196 0.8234 0.7232 0.8332 0.8162 0.8527 0.1241 0.5585 0.4461 0.5871 0.5537 0.6229
PBR $\mathbf{0.2267^{*}}$ $\mathbf{0.8592^{*}}$ $\mathbf{0.7780^{*}}$ $\mathbf{0.8754^{*}}$ $\mathbf{0.8640^{*}}$ $\mathbf{0.8902^{*}}$ $\mathbf{0.1408^{*}}$ $\mathbf{0.5800^{*}}$ $\mathbf{0.4463^{*}}$ $\mathbf{0.5949^{*}}$ $\mathbf{0.5847^{*}}$ $\mathbf{0.6424^{*}}$
[…]
Base 0.2267 0.8616 0.7900 0.8863 0.8926 0.9007 0.1217 0.5943 0.5322 0.6605 0.6444 0.6929
HyDE 0.2196 0.8759 0.7995 0.8974 0.8998 0.9044 0.1313 0.5943 0.5298 0.6538 0.6659 0.7010
Query2Term 0.2100 0.8544 0.7637 0.8739 0.8926 0.8989 0.1360 0.5776 0.5155 0.6290 0.6348 0.6754
MILL 0.2267 0.8663 0.7804 0.8819 0.7804 0.8680 0.1289 0.5967 0.5107 0.6362 0.5107 0.6278
CoT 0.2196 0.8425 0.7637 0.8707 0.8783 0.8918 0.1384 0.5609 0.4964 0.6148 0.5967 0.6570
ThinkQE 0.2124 0.8807 0.8108 0.9012 0.9021 0.9126 0.1384 0.6072 0.5080 0.6321 0.6445 0.7013
PBR $\mathbf{0.2315^{*}}$ $\mathbf{0.8902^{*}}$ $\mathbf{0.8138^{*}}$ $\mathbf{0.9072^{*}}$ $\mathbf{0.9045^{*}}$ $\mathbf{0.9191^{*}}$ $\mathbf{0.1408^{*}}$ $\mathbf{0.6205^{*}}$ $\mathbf{0.5489^{*}}$ $\mathbf{0.6612^{*}}$ $\mathbf{0.6874^{*}}$ $\mathbf{0.7060^{*}}$
Table 2: Retrieval performance in LongMemEval. The best results are in bold, and the second-best results are underlined. “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
4 Experiment
In this section, we conduct experiments to address the following research questions:
  • RQ1: Does the proposed PBR method improve retrieval performance on personalized corpus?
  • RQ2: What are the individual contributions of the P-PRF and P-Anchor modules to the overall performance?
  • RQ3: How does information propagation within the P-Anchor module affect the alignment between the query and user-specific corpora?
4.1 Experiment Setting
Datasets.
We evaluate our method on two recent benchmarks tailored for personalized retrieval. PersonaBench (Tan et al. 2025) simulates user-specific queries over synthetic private data (e.g., preferences, behaviors, demographics), requiring models to align responses with personalized context. It includes both explicit and implicit user traits across diverse domains. LongMemEval (Wu et al. 2025) targets long-term interactive memory, where each user is associated with multi-turn dialogue traces. We use two subsets: LongMemEval-s (sparse history) and LongMemEval-m (dense history), to evaluate the effectiveness of retrieval under both short- and long-horizon personalization. More details are provided in Appendix 1.
Evaluation Metrics.
We report performance using two standard retrieval metrics: Recall@K (R@K), which measures the proportion of relevant items ranked within the top-K retrieved results, and NDCG@K (N@K) to reflect retrieval quality.
Baselines.
We compare our method with six strong baselines: Base, a static query retriever without expansion; HyDE (Gao et al. 2023), which generates hypothetical answers to guide retrieval; Query2Term (Jagerman et al. 2023), which performs keyphrase-based query expansion; MILL (Jia et al. 2023), which leverages multi-level logical structures; CoT (Wei et al. 2022), which applies chain-of-thought reasoning for better ranking; and ThinkQE (Lei, Shen, and Yates 2025).
Implement details.
To verify the robustness of PBR, all methods are evaluated under three retrieval backbones: multi-qa-MiniLM-L6-cos-v1 (Wang et al. 2020), all-MiniLM-L6-v2 (Wang et al. 2020), and bge-base-en-v1.5 (Chen et al. 2024). For a fair comparison, all methods are using GPT-4o-mini (Achiam et al. 2023), and run 5 times. In PBR framework, we set $k_{1}=5$,$m=5$, $\theta=0.75$, and $k_{2}=10$ across Perosnabench and LongMemEval-s. The larger $k_{2}=50$ is particularly suited for LongMemEval-m, which contains richer user histories. More details are provided in Appendix 2.
4.2 Overall Performance (RQ1)
Figure 3: Overall performance in PersonaBench comparison across three retrieval models.
Figure 4: Parameter sensitivity analysis on P-Anchor of $\theta$ and $k_{2}$ settings.
To comprehensively evaluate the effectiveness of PBR, we conduct experiments on two benchmarks. PersonaBench contains 6 users, each with a dedicated corpus and about 50 user queries characterized by ambiguous intent and highly personalized language. In contrast, LongMemEval simulates a large-scale factual QA scenario with 500 distinct user queries, each paired with a relevant memory corpus. The former highlights soft personalization challenges (e.g., preferences and social traits), while the latter emphasizes hard factual grounding under long-context retrieval.
PBR resolving semantic ambiguity through user-specific expansion. Queries in PersonaBench often exhibit semantic openness and personal references—such as “Where is my hometown?” or “What is my favorite color?”—which cannot be resolved without grounding in the user’s unique corpus. As shown in Table 3.2, our PBR framework achieves the best overall performance across all retrievers (R@5: 0.4527, N@5: 0.3819), significantly surpassing the best baseline (ThinkQE: 0.4098 / 0.3484). This gain stems from PBR’s ability to generate user-style pseudo-feedback that anchors the query into a meaningful semantic space, reducing ambiguity and improving retrieval accuracy.
PBR excels in factual retrieval by boosting top-ranked precision. In contrast to PersonaBench, queries in LongMemEval are fact-oriented with specific personalization requirements, each associated with a specific memory corpus. In this setting, PBR boosts top-ranked retrieval performance. In Table 3.4, PBR also outperforms all baselines on LongMemEval, particularly on R@1, where it achieves 0.2315 on the -s version and 0.1408 on the -m version using bge. Compared to the best-performing baseline (0.2267 / 0.1384), PBR achieves absolute gain on top-1 precision. This suggests that our style simulation and structural anchoring more precisely locate relevant memory, elevating correct answers to top-1 even in long-context settings.
4.3 Ablation Study (RQ2)
To investigate the individual roles of P-PRF and P-Anchor in our framework, we perform a module-wise ablation across three retrievers in PersonaBench, as shown in Table 6. Each model variant is evaluated under both R@5 and N@5. More details are provided in Appendix 3.
PBR w/o P-Anchor w/o P-PRF
Metrics R@5 N@5 R@5 N@5 R@5 N@5
[…]
Overall 0.5035 0.4201 0.4871 0.4100 0.3457 0.2877
Basic information 0.5091 0.3647 0.4833 0.3570 0.3318 0.2242
Preference (hard) 0.4049 0.4175 0.4390 0.4295 0.2683 0.2981
Social 0.5541 0.4914 0.5164 0.4529 0.4277 0.3851
Preference (easy) 0.5321 0.5129 0.5192 0.5160 0.3590 0.3414
[…]
Overall 0.4516 0.3855 0.4382 0.3729 0.2860 0.2449
Basic information 0.4515 0.3485 0.4379 0.3393 0.1985 0.1617
Preference (hard) 0.4341 0.4352 0.4244 0.4201 0.2878 0.2718
Social 0.4494 0.3777 0.4333 0.3639 0.4179 0.3387
Preference (easy) 0.4840 0.4800 0.4712 0.4587 0.3846 0.3634
[…]
Overall 0.4029 0.3402 0.3921 0.3256 0.2699 0.2214
Basic information 0.4121 0.3057 0.3970 0.2958 0.2576 0.1804
Preference (hard) 0.3707 0.3889 0.3902 0.3778 0.2146 0.2173
Social 0.3657 0.3089 0.3431 0.2822 0.3261 0.2816
Preference (easy) 0.4904 0.4734 0.4744 0.4582 0.2949 0.2785
Average 0.4527 0.3819 0.4391 0.3695 0.3005 0.2513
Table 3: Ablation study of PBR in PersonaBench.
P-PRF drives major performance gains in personalized retrieval. Removing P-PRF leads to a substantial performance drop across all retrievers (e.g., on all-MiniLM, R@5 falls from 0.4516 to 0.2860), with the most severe decline observed on personalization-heavy subsets. This underscores the importance of P-PRF in generating semantically rich and stylistically aligned pseudo queries that effectively capture user intent.
Effectiveness of P-PRF Components. To assess how P-PRF enhances intent modeling, we perform ablation studies in PersonaBench and three retrievers (Figure 3). Removing either the roughly-aligned pseudo-utterance or logically-structured pseudo-reasoning consistently degrades performance in R@5 and N@5. The former harms lexical coverage, while the latter impairs goal-oriented reasoning, reducing ranking precision on inference-heavy datasets.
P-Anchor promotes semantic alignment in structured corpora. Removing P-Anchor leads to a consistent drop (e.g., N@5 from 0.3819 to 0.3695), especially in Basic Information and Social, where user corpora exhibit clearer structure. In contrast, in non-clustered tasks like Preference (hard), it offers limited gain and may overfit. P-Anchor is most effective when user context is semantically coherent.
4.4 Parameter Sensitivity Analysis (RQ3)
(a) Query-GT Comparison
(b) HyDE-GT Comparison
(c) PBR-GT Comparison (Ours)
Figure 5: Visualization of retrieval distance of the query in the Figure 1 compared to ground truth (GT) under different methods: (a) vanilla query, (b) HyDE, and (c) our proposed PBR.
To verify how information propagation within the P-Anchor module influences the alignment between queries and user-specific semantic space, we conduct a sensitivity analysis over two structural hyperparameters: the semantic threshold $\theta$ and the neighbor size $k_{2}$. The results are shown in Figure 4.
Moderate propagation improves alignment. When $\theta\in[0.65,0.75]$ and $k_{2}=5$, both R@5 and N@5 reach consistent peaks across datasets, indicating that controlled expansion of semantic neighborhoods facilitates better anchoring and personalization.
Over-propagation causes semantic drift. As $\theta$ and $k_{2}$ continue to increase, performance either saturates or slightly declines (e.g., at $\theta=0.8$, $k_{2}=10$), suggesting that excessive propagation introduces noise and weakens the semantic specificity of the alignment.
4.5 Visualization Case Study
To verify the practical effectiveness of our proposed PBR framework, we conduct a retrieval visualization case study. Figure 5 visualizes the t-SNE distribution of user queries, ground truths (GT), and generated queries under different methods. The results show that traditional expansion methods produce generic results misaligned with different users. Measuring cosine similarity to user-specific ground truth, PBR achieves higher separation between users (0.31 vs. 0.27), outperforming HyDE (0.04 vs. –0.01) and the original query (0.05 vs. 0.10). This highlights PBR’s effectiveness in embedding user-personalized semantics.
P-PRF enables semantically diverse and intent-aligned queries. Compared to vanilla queries or HyDE-generated expansions, PBR-generated queries are semantically closer to ground truth responses in both user clusters. The trajectory of PBR vectors more accurately captures user-relevant preferences, confirming its effectiveness in generating expressive and personalized query variants.
Improved alignment leads to more successful retrieval. Unlike HyDE, which generates a generic QE of all users, PBR yields multiple expansions (e.g., PBR user1 and PBR user2) that better cover the user-specific semantic space. This personalized QE allows the retriever to more precisely locate relevant results near each GT, reducing retrieval distance and improving semantic alignment.
5 Literature Review
5.1 Query Expansion in RAG System
Query expansion (QE) improves retrieval performance by simulating PRF. With the rise of LLMs, recent studies have explored LLMs as zero-shot QE engines. HyDE (Gao et al. 2023) and Query2Doc (Wang, Yang, and Wei 2023) demonstrate that prompting LLMs can yield precise expansions with PRF. Other works enhance semantic diversity or reasoning depth: MILL (Jia et al. 2023) employ mutual verification, and GRF (Mackie, Chatterjee, and Dalton 2023) apply LLM-based relevance feedback. ThinkQE (Lei, Shen, and Yates 2025) models expansion as a reasoning process, while LLMlingua (Jiang et al. 2023) focuses on compressing expansion prompts. GAR (Xia et al. 2024) further introduces a knowledge graph-aware QE framework.
However, existing work overlooks personalization and fails to align QE with user-specific semantic spaces. In contrast, our PBR generates style-aware and structure-aware expansions, bridging this gap for personalized retrieval.
5.2 Personalized RAG System
Personalized RAG system has recently emerged as a crucial direction for adapting LLMs to individual users. A recent survey by Li et al. (Li et al. 2025, 2023) provides a comprehensive overview, highlighting the importance of personalization in RAG. Approaches such as UniMS-RAG (Wang et al. 2024) integrate multi-source signals for dialogue personalization, while PersonaRAG (Zerhoudi and Granitzer 2024) enhances personalization via role-based agent frameworks. Salemi et al. (Salemi et al. 2024) further investigate LLM personalization by modeling language preferences and user goals. Complementary work like Wu et al. (Wu et al. 2024) explores the role of explicit user profiles, and Cohn et al. (Cohn et al. 2025) leverage interaction logs for educational agent adaptation. Domain-specific applications, such as personalized care systems (Yang et al. 2025) and memory-based assistant evaluation (Wu et al. 2025; Xu et al. 2025b, a), highlight the importance of long-term memory and evolving user states in effective RAG.
Despite recent advances, little attention has been given to improving retrieval accuracy at the pre-retrieval stage. PBR enables fine-grained personalization by jointly modeling the user’s style and corpus structure before retrieval.
6 Conclusion
This paper introduces PBR, a Personalize-Before-Retrieve framework for personalized retrieval. PBR expands user query prior to retrieval by integrating two components: P-PRF, a style-aware pseudo query generator, and P-Anchor, a graph-based corpus alignment module. Our results highlight the importance of pre-retrieval personalization for improving user-centric retrieval performance. Future work will investigate generalizing PBR to broader information-seeking tasks beyond query-based retrieval.
Appendix 1: Datasets Details
PersonaBench.
PersonaBench (Tan et al. 2025) [^1] https://github.com/SalesforceAIResearch/personabench is a benchmark dataset designed to evaluate personalized retrieval grounded in user-specific contexts. It includes six users, each associated with a heterogeneous personal corpus composed of (i) conversations with friends (Conv.), (ii) dialogues with AI assistants (AI.), and (iii) e-commerce purchase histories (E-com.). For each user, 40–50 short and ambiguous natural language queries are provided (e.g., “Do you know what my favorite color is?”), which require accurate grounding in the user’s prior behavior. This setup enables evaluation of models’ ability to align with diverse personal semantics. Table 4 summarizes the query and corpus statistics across users.
User Queries Corpus Conv. AI E-com.
1 48 110 84 23 3
2 43 90 78 8 4
3 42 64 51 12 1
4 46 85 71 14 0
5 44 84 59 21 4
6 40 94 79 14 1
Sum 263 527 422 92 13
Avg 43.83 87.83 70.33 15.33 2.17
Table 4: Statistics of the PersonaBench dataset across six users. The Corpus is the sum of Conv., AI., and E-com..
LongMemEval.
LongMemEval (Wu et al. 2025) [^2] https://github.com/xiaowu0162/LongMemEval is a benchmark designed to evaluate the long-term memory capabilities of large language models, particularly their ability to retrieve task-relevant information from extended user-specific corpora. Each instance in the dataset consists of a factual question paired with a synthetic memory corpus simulating historical user data. The dataset includes two settings: LongMemEval-s, where each question is associated with approximately 50 memories, and LongMemEval-m, where each corpus contains over 500 memories. In total, both settings consist of 500 questions, allowing evaluation of memory retrieval precision under varying context lengths. This benchmark enables controlled analysis of personalized retrieval performance under realistic long-context scenarios.
Statistic LME-s LME-m
Total Questions 500 500
Total Memory 25,112 250,948
Min Memory per Question 39 501
Max Memory per Question 66 506
Avg Memory per Question 50.22 501.90
Table 5: Statistics of the LongMemEval dataset (LME). Both LME-s and LME-m contain 500 questions, with different sizes of associated memory corpora.
Appendix 2: Implement Details
More Detailed Implement
Base Settings
: All retrieval experiments are performed on a single NVIDIA A100 80GB GPU under Ubuntu 22.04. Query expansion is generated using GPT-4o-Mini with a temperature of 0. And each experiment we run 5 times.
Retrieval
: To verify the robustness of our approach, all methods are evaluated under three retrieval backbones: multi-qa-MiniLM-L6-cos-v1333https://huggingface.co/sentence-transformers/multi-qa-MiniLM-L6-cos-v1, all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, and bge-base-en-v1.5555https://huggingface.co/BAAI/bge-base-en-v1.5. And FAISS to store the embedding vectors for efficient retrieval. For retrieval, we use Euclidean distance.
Prompt Templates for Query Expansion Methods
Prompt of Baselines
To ensure a fair comparison, we implemented the prompt-based query expansion strategies of representative baseline methods. These prompts were carefully designed to align with the original paper.
HyDE Prompt Please write a paragraph that answers the question. Question: {query}
Output:
Query2Term Prompt Answer the following question: {query}
Give the rationale before answering.
MILL Prompt What sub-queries should be searched to answer the following query? Please generate 5 sub-queries with their related passages. Question: {query}
You should only return a python list like:
["query1 passage1", "query2 passage2", ..., "query5 passage5"]
(no comments, no markdown) without any other words and explanation.
Chain-of-Thought Prompt Solve the question step-by-step. Question: {query}
ThinkQE Prompt Given a question ’q’ and its possible answering passages (most are wrong): 1. {d1}; 2. {d2}; 3. {d3}; 4. {d4}; 5. {d5}
Please write a correct answering passage.
Use your own knowledge, not just the example passages!
Prompt Design for Our Method
Our personalized query expansion framework (PBR) involves a two-stage prompt design to jointly simulate user-specific expression (P-PRF) and adaptively anchor reasoning paths (P-Anchor). Below we present the two corresponding prompts:
P-PRF-roughly You are to generate 10 natural candidate utterances the user might say, inspired by the dialogue history and the current question. Context User dialogue history (for style imitation): {history} Current question (to inspire the utterances): {query} Guidelines 1. Generate 10 fluent, natural utterances the user might plausibly say. 2. Do NOT just paraphrase; include variations in tone, emphasis, or context. 3. Each utterance should exceed 25 words. 4. Reflect the style and tone consistent with the document. 5. Return ONLY valid JSON in this format (no comments, no markdown): { "candidates": ["...", "...", "..."] }
P-PRF-logically Solve the question step-by-step, inspired by the user dialogue history. Context User dialogue history (for style imitation): {history} Current question (to inspire the reasoning): {query} Output: step-by-step explanation that maintains user-specific tone and reasoning patterns.
These two prompts work in tandem: P-PRF synthesizes diverse user-style-aligned utterances, while P-Anchor explicitly grounds these expansions via structured reasoning. Together, they allow the model to generate personalized and semantically anchored queries before retrieval.
Appendix 3: Extend Experiment
Time Efficiency
Table 6 reports the latency of each stage in the ThinkQE and PBR methods, measured in seconds. Both methods share similar preprocessing overheads during index construction and user information retrieval. PBR introduces an additional graph-building step (0.0008 s) to enhance personalized query expansion, which incurs negligible computational cost. The dominant latency for both methods arises from the query expansion generation stage, accounting for approximately 2.5–2.7 s, while the final retrieval stage remains lightweight. Overall, PBR achieves personalization with marginal additional latency about 0.18 s compared to ThinkQE.
Stage ThinkQE PBR
1. Build index 0.0797 0.0797
2. Build graph 0.0008
3. Retrieval user info 0.0043 0.0043
4. Call API gen QE 2.4999 2.6646
5. Final retrieval 0.0064 0.0167
Table 6: Time latency of each stage in ThinkQE and PBR.
Senitivity of $k_{1}$
We further conducted a hyperparameter sensitivity analysis on the LongMemEval-S dataset, focusing on the impact of $k_{1}$, which controls the number of top-ranked candidates used in the retrieval of the P-PRF module. As shown in Table 7, increasing $k_{1}$ from 3 to 10 consistently improves both recall and ndcg across all evaluation cutoffs. The gains are particularly evident at higher recall thresholds (e.g., R@5 increases from 0.8253 to 0.8401), indicating that a larger candidate pool allows the model to capture more diverse yet relevant context signals. We set $k_{1}=5$ as the default in the main experiments.
Table 7: Hyperparameter sensitivity on the LongMemEval-S dataset under different values of $k_{1}$.
$k_{1}$ R@1 N@1 R@3 N@3 R@5 N@5
3 0.2076 0.8258 0.7227 0.8402 0.8253 0.8519
5 0.2100 0.8210 0.7279 0.8444 0.8282 0.8598
10 0.2153 0.8305 0.7470 0.8504 0.8401 0.8661
Generation Performance
To further evaluate the effect of personalized query expansion on generative outcomes, we conducted an additional experiment on the LongMemEval-S dataset under multi-qa-MiniLM-L6-cos-v1 retriever. Following prior work on RAG evaluation (Adlakha et al. 2024), we adopt a subspan Exact Match (EM) metric to account for partial correctness in generative responses. As shown in Table 8, PBR outperforms baseline query expansion approaches such as HyDE and CoT. Notably, PBR achieves the highest EM score of 42.6%, surpassing the non-personalized CoT (41.0%) and ThinkQE (41.8%). These results demonstrate that integrating personalized query representation not only enhances retrieval accuracy but also yields measurable gains in final generation quality, confirming the cross-stage benefit of the proposed pre-retrieval personalization.
Table 8: Generation performance (EM) on the LongMemEval-S dataset.
Method EM (%)
Base 40.0
HyDE 41.4
Query2Term 40.2
MILL 40.8
CoT 41.0
ThinkQE 41.8
PBR 42.6
PBR w/o P-Anchor w/o P-PRF w/o both
Metrics R@5 N@5 R@5 N@5 R@5 N@5 R@5 N@5
[…]
Overall 0.5035 0.4201 0.4871 0.4100 0.3457 0.2877 0.4484 0.3669
Basic information 0.5091 0.3647 0.4833 0.3570 0.3318 0.2242 0.4515 0.3088
Preference (hard) 0.4049 0.4175 0.4390 0.4295 0.2683 0.2981 0.3659 0.3759
Social 0.5541 0.4914 0.5164 0.4529 0.4277 0.3851 0.4852 0.4356
Preference (easy) 0.5321 0.5129 0.5192 0.5160 0.3590 0.3414 0.4904 0.4582
[…]
Overall 0.4516 0.3855 0.4382 0.3729 0.2860 0.2449 0.3783 0.3074
Basic information 0.4515 0.3485 0.4379 0.3393 0.1985 0.1617 0.3515 0.2644
Preference (hard) 0.4341 0.4352 0.4244 0.4201 0.2878 0.2718 0.4000 0.3547
Social 0.4494 0.3777 0.4333 0.3639 0.4179 0.3387 0.3921 0.3048
Preference (easy) 0.4840 0.4800 0.4712 0.4587 0.3846 0.3634 0.4295 0.4199
[…]
Overall 0.4029 0.3402 0.3921 0.3256 0.2699 0.2214 0.3738 0.3015
Basic information 0.4121 0.3057 0.3970 0.2958 0.2576 0.1804 0.3970 0.2748
Preference (hard) 0.3707 0.3889 0.3902 0.3778 0.2146 0.2173 0.3268 0.3343
Social 0.3657 0.3089 0.3431 0.2822 0.3261 0.2816 0.3204 0.2799
Preference (easy) 0.4904 0.4734 0.4744 0.4582 0.2949 0.2785 0.4583 0.4065
Average 0.4527 0.3819 0.4391 0.3695 0.3005 0.2513 0.4002 0.3253
Table 9: Ablation study of PBR across different retrievers and subtasks.
Dataset LongMenEval-s LongMenEval-m
Metrics R@1 N@1 R@3 N@3 R@5 N@5 R@1 N@1 R@3 N@3 R@5 N@5
[…]
PBR 0.2100 0.8210 0.7279 0.8444 0.8282 0.8598 0.1217 0.5537 0.4320 0.5776 0.5537 0.6177
w/o P-Anchor 0.2100 0.8129 0.7199 0.8385 0.8239 0.8420 0.1169 0.5456 0.4268 0.5623 0.5480 0.6127
w/o P-PRF 0.1838 0.7566 0.6277 0.7540 0.7208 0.7739 0.1122 0.4893 0.3484 0.4961 0.4368 0.5303
w/o Both 0.1885 0.7924 0.6730 0.8020 0.8067 0.8301 0.1098 0.5251 0.4057 0.5559 0.5394 0.6038
[…]
PBR 0.2267 0.8592 0.7780 0.8754 0.8640 0.8902 0.1408 0.5800 0.4463 0.5949 0.5847 0.6424
w/o P-Anchor 0.2191 0.8522 0.7628 0.8622 0.8511 0.8860 0.1360 0.5795 0.4435 0.5915 0.5819 0.6356
w/o P-PRF 0.1957 0.7566 0.6563 0.7765 0.7589 0.8017 0.1217 0.4797 0.3628 0.4961 0.4678 0.5415
w/o Both 0.2196 0.8234 0.7399 0.8477 0.8568 0.8731 0.1265 0.5322 0.4415 0.5857 0.5346 0.6191
[…]
PBR 0.2315 0.8902 0.8138 0.9072 0.9045 0.9191 0.1408 0.6205 0.5489 0.6612 0.6874 0.7060
w/o P-Anchor 0.2239 0.8850 0.8121 0.9037 0.8993 0.9154 0.1304 0.6125 0.5322 0.6579 0.6721 0.7007
w/o P-PRF 0.2029 0.8234 0.7446 0.8517 0.8449 0.8704 0.1002 0.5298 0.4344 0.5698 0.5680 0.6125
w/o Both 0.2267 0.8616 0.7900 0.8863 0.8926 0.9007 0.1217 0.5943 0.5322 0.6605 0.6444 0.6929
Table 10: Ablation results of PBR on LongMenEval-s and LongMenEval-m datasets.
Ablation Study in Two Datasets
Complementary Ablation Insights.
Across both the retrieval backbones and datasets, we observe a consistent performance drop when either component (P-Anchor or P-PRF) is removed, confirming their complementary contributions to the overall effectiveness of the PBR framework.
First, removing P-PRF leads to the most significant degradation, especially on more complex queries such as Preference (hard) and datasets like LongMenEval-m, where user intent is more nuanced. For example, on multi-qa-MiniLM in Table 6, R@5 drops from 0.5035 to 0.3457 without P-PRF, and similarly in Table 6, LongMenEval-s sees a drop in R@3 from 0.7780 to 0.6563 without P-PRF. This highlights the critical role of user-aware pseudo relevance feedback in capturing latent intent.
Second, removing P-Anchor has a more moderate but consistent impact, especially in tasks requiring robust grounding of queries within diverse semantic corpora (e.g., Social and Preference (easy)). This suggests that style-consistent anchoring helps stabilize query representations, particularly when user content is lexically varied.
Finally, the combination of removing both components consistently results in a further drop, but not always the worst performance. This indicates a certain degree of redundancy or shared effect between P-Anchor and P-PRF, suggesting that while they operate on different principles (style anchoring vs. semantic feedback), they jointly reinforce personalized query understanding. On average (last row, Table 6), the full PBR outperforms the no-component baseline by +13.1% R@5 and +17.4% N@5.
Effectiveness of P-PRF in Two Datasets
Figure 6: Overall performance in Personabench comparison across three ablation retrieval modes.
Visualization-based Analysis.
Figure 6 presents a comparative bar plot of PBR and its ablated variants across three retrievers (multi-qa-MiniLM, all-MiniLM, bge-base-en) and three evaluation datasets (PersonaBench, LongMenEval-s, LongMenEval-m). Each subplot reports the R@5 and N@5 metrics, offering insight into both retrieval recall and ranking quality.
(1) On PersonaBench, PBR consistently outperforms all variants across all retrievers. Notably, the drop in N@5 when removing P-PRF is more severe than the drop in R@5, indicating that pseudo-relevance feedback contributes more to fine-grained ranking alignment than coarse recall.
(2) On LongMenEval-s, where queries are more succinct, the gap between PBR and its variants becomes more pronounced, especially under bge-base-en. This suggests that PRF-based refinements are crucial for handling underspecified queries by injecting user-grounded semantics into the expansion.
(3) On LongMenEval-m, the performance degradation from removing P-PRF is still significant but relatively less than in LongMenEval-s. This is likely due to longer query inputs providing stronger initial signals, slightly compensating for the lack of PRF. However, the logical and rough removal of PRF still lags behind full PBR, validating the importance of structurally grounded expansion paths.
Overall, the visual analysis confirms the complementary roles of anchor-based grounding and pseudo-feedback refinement in boosting personalized retrieval quality across varied retriever and dataset settings.